### v1

# AccidentBlip2: Accident Detection With Multi-View MotionBlip2

Abstract
Multimodal Large Language Models (MLLMs) have shown outstanding capabilities in many areas of multimodal reasoning. Therefore, we use the reasoning ability of Multimodal Large Language Models for environment description and scene understanding in complex transportation environments. In this paper, we propose AccidentBlip2, a multimodal large language model that can predict in real time whether an accident risk will occur. Our approach involves feature extraction based on the temporal scene of the six-view surround view graphs and temporal inference using the temporal blip framework through the vision transformer. We then input the generated temporal token into the MLLMs for inference to determine whether an accident will occur or not. Since AccidentBlip2 does not rely on any BEV images and LiDAR, the number of inference parameters and the inference cost of MLLMs can be significantly reduced, and it also does not incur a large training overhead during training. AccidentBlip2 outperforms existing solutions on the DeepAccident dataset and can also provide a reference solution for end-to-end automated driving accident prediction. The code will be released at: https://github.com/YihuaJerry/AccidentBlip2.git
I INTRODUCTION
Multimodal Large Language Models (MLLMs)[1] that can detect and accurately determine the occurrence of accidents are important to the safety field of autonomous driving. There is some precedent for people to detect about vehicle behaviour. Normally, the vehicle will be driven according to its surroundings and in accordance with the traffic rules. In addition, in complex traffic systems, vehicles can stop, change lanes or even reverse. Therefore we include all these phenomena in the vehicle surroundings. Although these phenomena are all included in the context of vehicle environment sensing, the inherent complexity of the traffic system will lead to difficulties in modeling road sensing. Therefore using traditional neural network perception methods often results in negative judgments.
((a)) Hints for hazardous situations
((b)) Tips for the security situation
Figure 1: Agent for accident prediction and road condition alerts: In response to the temporal pictures, 1(a) is the agent alerts the driver if there is a potential hazard or impending hazard in the sequence, 1(b) is if there are no safety hazards in the sequence pictures, the agent alerts the driver to drive safely.
However methods for purely visual accident situation detection for complex scenes are still lacking. But the ability of multimodal large language models to understand complex scenes makes them stand out in complex environments for autonomous driving tasks. Existing representative work typically uses multimodal bigram models to detect vehicles, people, etc. in the environment, thus enhancing reliability in autonomous driving. However, in complex traffic scenarios, accidents occur frequently, so the in-vehicle multimodal large language model can use its scenario understanding ability to determine the surrounding accident information for the self-driving vehicle and help the traffic manager to know whether there is any danger or collision in the surrounding environment.
In this paper, we propose AccidentBlip2, a multimodal large language model for accident judgment for time-series scenarios. We aim to advance the practical application of multimodal large models in complex traffic scenarios. With multimodal large models, we implemented a framework for compiling 6-view camera images into a multimodal input temporal Qformer with vision transformers. Subsequently, these temporal tokens generated by the vision transformer can be used to train the temporal Qformer. In practice, the token of the temporal images is passed from the Qformer to the visual encoder of the multimodal large language models, whereby the reasoning power of the multimodal large language models is utilized to predict and determine whether an accident has occurred. Multimodal large language models can also interact with the driver through language to more accurately sense the environment on the road and detect any hazards in the way.
In addition to the perceptual analysis of the surrounding scene of a single vehicle, we have also developed a multi vehicle collaborative end-to-end scenario perception scheme to compensate for some of the blind spots and deficiencies of single-vehicle perception. We extended the single-vehicle environmental experiments to end-to-end driving scenarios, testing the accuracy of accident judgment and perception in addition to self-vehicle and multi-vehicle association, respectively. Overall, our main contributions are as follows:
  • 1)
    We propose a novel visual traffic accident judgement agent, which addresses the prediction of accident-specific occurences and alerts drivers to the possibility of an accident in complex driving environments.
  • 2)
    We present a framework for end-to-end accident prediction based on multimodal large language models, enabling the large models to determine whether an accident has occurred or there is a risk of an accident in or around a multi-vehicle system based on end-side vehicle information.
II RELATED WORK
II-A MultiModal Large Language Model
With the advent of GPT4, a large number of large language models began to explore multimodal capabilities. These large language models use multimodal information to fine-tune the language model to enhance the large language models’ fusion ability and comprehension in different modalities. Among them, the visual-verbal multimodal large language models, led by GPT-4V and Llava-v1.5, have spawned multimodal large language models models applicable to different scenarios. In addition owlViT, Qwen-VL, and other multimodal large language models for target detection are fine-tuned for specific scenes by combining specific datasets, allowing them to accomplish specific visual tasks with user-prompt commands. For temporal multimodal information, researchers have also introduced Video-Llava and Video-Llava, which incorporate the processing capabilities of temporal pictures and videos into a large language model. There are also researchers currently fine-tuning data from audio modalities for large language models, such as Qwen-Audio, which facilitates the development of large language models for audio modalities. These cross-modal fusion of large language models facilitated by large language models in different scenarios provide quality solutions for different tasks in the open-world.
II-B MLLMs for Automatic Driving
Recently, with the rapid development of autonomous driving models, big language models are gradually appearing in the field of intelligent driving and Telematics. 2023 will see the emergence of uniad, which completely applies big language models to the field of autonomous driving. for the first time, uniad integrates the three important characters of autonomous driving: perception, decision-making, and planning in a unified network architecture, effectively reducing the loss of information between different independent modules. Also, researchers developed a multimodal large language model that can focus on driver intent called CAVG. CAVG consists of five specialized encoders: a text encoder, an emotion encoder, a visual encoder, a context encoder, and a cross-modal encoder. These encoders, in combination with their corresponding decoders, allow the large language model to handle different modalities of autonomous driving tasks. In the perception task, DRIVEGPT4 uses YOLOv8 to detect common targets such as vehicles in each frame of the video and sends the obtained border coordinates to ChatGPT as linguistic information. However, since DriveGPT4 only does environment-aware tasks for a single vehicle, it cannot be practically deployed for end-to-end tasks with multi-vehicle linkage.
Figure 2: Architecture of MotionBlip2: We propose a multi-view temporal prediction scheme. We firstly feed the initialized query $Q_{0}$ into the Qformer corresponding to the first frame of the multi-view images. , and at the same time input the features of the first frame of the multiview processed by ViT into the first Qformer. A new query $Q_{1}$ will be generated by the Qformer inference and input into the next Qformer along with the features of the ring view map of the current frame processed by ViT. After completing the last frame of image processing, we input the generated query $Q_{n}$ into the MLP and perform autoregressive computation, if there is a risk of accident we will alert the driver and vice versa we will remind the driver to drive with caution.
II-C Accident Judgment
Judgment of traffic accidents as the most concerned research area in autonomous driving safety, many researchers have done a lot of work on accident Judgment. Based on traditional detection methods, the researchers used the front view of the vehicle in conjunction with a neural network for temporal prediction, such as an LSTM or RNN, to warn drivers of a crash[10]. However, this method can only sense the vehicle travelling status around the bicycle, and cannot sense the danger in the complex traffic environment. The probability of correct accident prediction decreases if complex weather is encountered. With the rapid development of large language models, there have been researchers using large language models in accident perception, such as AccidentGPT. however, due to the fact that its corpus is directly retrieved through large language models, it is not able to reason about the perception of brand new environments in complex situations.
III METHODOLOGY
Blip2 is not able to directly process the 6-view pictures inputs, nor is it able to directly use temporal 6-view pictures inference. In this section, we use the CARLA[11] simulator’s simulation dataset and vision transform to enhance Blip2’s ability to solve temporal 6-view pictures. Thus we hope that the trained motion Blip2 will be able to reason over the temporal 6-view images that can be used for inference. Notably, our framework can be extended to multi-vehicle end-to-end autonomous driving simulation scenarios.
III-A Multi-view Input and Temporal Inference
We introduce a perceptual framework that leverages multimodal large language models, comprising of two key components: multi-view image processing and temporal inference. In the multi-view processing stage, we utilize ViT-14g to handle multiple views and extract the relevant image features. These features capture important visual information from different perspectives. Moving to the temporal inference stage, we employ queries as carriers of temporal information for forward transfer. These queries encapsulate contextual and temporal cues that facilitate understanding and reasoning over time. By combining these two stages, our framework enables comprehensive perception and inference in a multimodal manner.
Our visual encoder uses ViT-14g from EVA-CLIP[13]. Vision Transformer first reads the multiple-view images provided by the camera and extracts features separately after resizing each image. We use the pre-trained vision transformer as a backbone to extract the features $f_{t}$ the multi-view of $t$ frames, and input the features of each multi-view frame into the linear layer of the Qformer to do the inference. Before inputting the multi-view features of the $t$ frame into Qformer, the vision transformer will first spread the features of each image into one dimension, then stitch the features of the six images, and input the stitched features $f_{t}$ into Qformer’s cross-attention. At the same time, the learnable query $Q_{t}$ is input into only Qformer’s self-attention layer, and the new query is output through the feed forward layer.
Figure 3: Architecture of Motion Qformer: We replace the self-attention layer in Qformer with temporal attention layer, and encode the temporal information by inputting the query $Q_{n-1}$ generated in the previous frame to the attention head of temporal attention. The encoded token is input into cross attention and reasoned with the multi-view images feature $f$ generated in the current frame to generate the query $Q_{n}$ corresponding to the current frame by feed forward layer.
Depending on the type of data in autopilot, we need to propose a framework that can handle temporal images. We propose temporal self-attention, which inputs the multi-view images’ information for each frame through the temporal query $Q_{n}$, enter $Q_{n}$ into the corresponding Qformer respectively, and obtain the query $Q_{n+1}$ corresponding to the input of the next frame. Taking advantages of attention mechanism in Qformer, it interacts with extracted features from last output of Qformer, as shown in 4. Each frame’s muti-view features $f_{t}$ present in attention with temporal query $Q_{n}$, written in 1,
$Attn(Q_{n-1},K_{Car},V_{Car})=Softmax(\frac{Q_{n}\cdot K_{Car}^{T}}{\sqrt{d_{k}}})V$ (1)
where $Q_{n-1}\in\mathbb{R}^{N_{Q}\times D}$ denotes the query feature from last timestamp, which incurs to generate the output of current state $Q_{n}$. After that, in order to present the current image view. the current state $Q_{n}$ combines the image features from ViT-g, denoted as $f_{t}$. This cross attention mechanism can be described as
$Q=Q_{n}W^{Q}\hskip 10.00002ptK=V=f_{t}W^{K}$ (2)
$CrossAttn(Q_{n},f_{t})=Softmax\left(\frac{QK^{T}}{\sqrt{d_{2}}}\right)V$ (3)
where $Q$ denotes the projection layer from the feature $Q_{n}$, $K$ and $V$ denotes the features from $f_{t}$.
III-B Multi-Vehicle End-to-End Sensing
To ensure reliable vehicle environment perception, relying solely on a single vehicle can be an inadequate approach due to the existence of blind zones in visual perception. To address this challenge, we have implemented a vision-only perception scheme for multi-vehicle interaction. In complex traffic systems, our vehicle neural network is designed not only to assess potential accidents involving the self-vehicle but also to evaluate the risk of accidents involving other vehicles. This is achieved through a multi-vehicle joint training method. Consequently, we have successfully trained a multi-vehicle accident-aware agent using data from the DeepAccident dataset.
Figure 4: Architecture of End-to-end AccidentBlip2: Our proposed solution to address multi-vehicle collaborative incident awareness. We follow the model architecture of MotionBlip2 for each vehicle.
In our multi-vehicle system, we employ a pre-trained ViT-14g model on each vehicle, denoted as $Car_{n}$, to extract the features $f$ from the current frame’s multi-view images of the self-vehicle. These features are then fed into our specially designed MotionBlip2 module. Additionally, for the queries generated by AccidentBlip2, we combine them with the multi-vehicle queries and input them into the Motion Qformer in the subsequent time step. Finally, we transform this high-dimensional query into a one-dimensional tensor to be used as input to the MLP (Multi-Layer Perceptron), as illustrated in Equation 4.
$\mathbf{X_{i}}=\text{MLP}(\text{concat}(Q_{1},Q_{2},Q_{3},Q_{4},...\text{dim}=0))$ (4)
We input the token output from the MLP at frame $i$ into embedding, and the embedded token $X_{i}$ is subjected to autoregression, from which we determine whether there is an accident between multi-vehicle systems. Each query $Q_{n}$ denotes one perspective from one vehicle $Car_{n}$ in multi-vehicle system, concatenated with one and the other, aggregating to be 24 viewpoints for the coordination of features.
IV EXPERIMENT
In this section, we validate our system on the simulation dataset DeepAccident. We test it separately for single-vehicle perception to multi-vehicle perception. More precisely, we will focus on explaining the following:
  • 1)
    Can AccidentBlip2 determine accidents in complex traffic systems more efficiently than other vision-only solutions used in time-series systems?
  • 2)
    Can AccidentBlip2 sense accidents occurring around it more accurately than a single-vehicle system in a multi-vehicle end-to-end system?
IV-A Training Data
For training and evaluation data, we have adopted a public open-source dataset, DeepAccident[2], taking advantanges of its full visual information upon vehicles and infrastructures. In the DeepAccident dataset, it consists of couples of scenarios that emulates those real-world potential car collisions, attached with 4 vehicles, 1 infrastructure and BEV point cloud annotations. Nevertheless, in contrast to those multimodal techniques, AccidentBlip2 is based on vision-exclusive paradigms, hence involving only visual images sampled from objects.
As proposed by DeepAccident[2], it consists of 12 types of accident scenarios that happens upon signalized and unsignalized intersections. The dataset has 57k annotated V2X frames, split by the ratio of 0.7, 0.15 and 0.15 for training, validation and test set respectively. Our training can be fit in a 4-GPU A6000 machine, which merely requires less than 24 hours for the result.
IV-B Baselines
Our main baseline is the common Video Large Language Model (Video-Llama). The use of Video-Llama is often seen when dealing with video tasks. we also compare it with Video-Vicuna, a fine-tuned version of Video-Llama. In addition to the Video-Llama family of video large language models, we also used Video-Llava, a video inference large language model of alignment before projection, as a comparison. Purely visual fine-tuned large language models like Llava and GPT-4V do not easily scale to temporal data inputs, and they are not superior to Video-Llama, so they are not included in the comparison.
IV-C Implementation Details
Our proposed model consists of ViT-14g as the visual encoder, Motion Qformer leveraged as extracting and concatenating features from both texts and images. While taking into account the optimum experiment for accident prediction, we train AccidentBlip2 with image size of $224\times 224$.
In our setup, the parameters of our model are frozen for the visual model ViT-14g and the language model OPT-2.7B[14], which have been initialized with pre-trained weights, equivalent as Blip-2[21]. Motion Qformer, as it sides with Qformer features, aims to fuse multiple view with collective vehicles together, loading and trained with tasks that supports for both prediction and Language Model outputs. Image input are passed through ViT layer into Motion Qformer in order to capture the temporal dynamic features within different frames.
During training, we selected a learning rate warm up for the initial 3 epochs, coupled with the Adam optimizer[1] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. The AccidentBlip2 model was trained for 8 epochs, using a cosine learning rate decay as the scheduler of learning rate up to $51e-$ 5. The training epoch is 6 and batch size is 8, varing from tasks involved with different numbers of vehicles. In terms of multiple scenarios, with sampling from timestamp 0 to T(final timestamp), the input shape is $X_{S}\in\mathbb{R}^{T\times N_{V}\times V_{C}\times C\times H\times W}$, while $T$ denotes the time length, $N_{V}$ represents the number of vehicle, $V_{C}$ denotes 6 separate view of the vehicle, encompassing Front, FrontLeft, FrontRight, Back, BackLeft and BackRight cameras, and $H$ and $W$ stand for the size of input image. Motion Qformer recursively interacts with the query output $Q_{T-1}\in\mathbb{R}^{N_{V}\times N_{Q}\times D}$, which spawns from the last timestamp viewpoints. It takes in and out the temporal multi-view features repetitively until reaching to its end, then LLM module is in charge of collecting the final Query Output $Q_{T}$ aggregating a comprehensive image embedding. Our training loss $Loss(p_{t})$, shown in following 5, is calculated based on Focal Loss, where $p_{t}$ denotes the probability of true class, $\alpha$ is used to balance the weights of positive and negative samples and $\gamma$ take controls the rate of samples detected.
$Loss(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ (5)
Owing to mitigating imbalance categories as our model predicts on accidents, $\alpha$ is set to be 0.25, while $\gamma$ lies to its default value 2.0, conducive to train the model gaining greater focus towards misclassified samples.
IV-D Evaluation
Table I presents the results of performances of different different V2X configuration inputs in our model.By designing different V2X configuration models to demonstrating the model’s performance, we have evaluated AccidentBlip2 in single-car scenario, multi-car scenario, including ego vehicle, other vehicle and their cars behind, along with full-view scenario by introducing infrastructure view. Benefiting from the six camera sensors installed across four vehicles, our Motion Qformer effectively captures the temporal features of multiple vehicles.
TABLE I: Different V2X Configuration Models
Ego Vehicle Other Vehicle Vehicles Behind Infrastructure Acc.
$\checkmark$ -- -- -- 66.5
$\checkmark$ $\checkmark$ -- -- 68.5
$\checkmark$ $\checkmark$ -- $\checkmark$ 66.7
$\checkmark$ $\checkmark$ $\checkmark$ -- 73.1
This results in an accuracy improvement of 2% compared to the single-vehicle scenario. Moreover, an apparent increment by 6.6% can be seen with the advent of multiple views, which shows that the model effectively takes into account the overview of accidents.
Our proposed model outperform baseline models in both single-vehicle and multi-vehicle configurations. Compared with these video large language models, shown in table II, our AccidentBlip2 achieves 66.5% with merely single car input. In comparison with V2XFormer, proposed with three different V2X-agent fusion setting, a significant increment can be observed in our model, with about 3% improvement in accuracy.
TABLE II: Performance Comparison Amongst Different Video Varge Language Models.
Model Data Sources Accuracy
Video-LLaMA single vehicle 51.34
multiple vehicles 33.33
Video-Vicuna single vehicles 36.43
multiple vehicles 31.87
Video-LLaVA single vehicle 49.7
multiple vehicles 47.7
V2XFormer single vehicle 43.8
4 vehicles + infra 69.5
AccidentBlip2 (ours) single vehicle 66.5
multiple vehicles 73.1
Our approach is designed to address the challenge of gradient shifts during training, which can lead to gradient explosion in the initial epochs. To mitigate this issue, we employ a warmup scheduler that makes slight adjustments to the learning rate within a minimal range. This helps stabilize the training process. Our multi-vehicle paradigm operates on four separate GPUs, allowing us to gather query results from each motion Qformer. Additionally, we conducted experiments to explore the impact of the initial weight of the Qformer, comparing pre-trained weights from Blip2 with default initialization. The results revealed a decrease in the overall performance of our language model when evaluated on the DeepAccident dataset. To overcome this, we utilize pre-trained weights from Blip2 to initialize AccidentBlip2, improving the model’s effectiveness in accident detection tasks.
V CONCLUSIONS
In this research paper, we propose a Motion Qformer-based accident detection framework named AccidentBlip2, which solely relies on vision inputs for analyzing road information. This framework introduces temporal attention in Blip2 by replacing self-attention mechanisms. It utilizes queries as carriers of temporal features, encoding the information from each frame into the temporal attention of the subsequent frame. This enables autoregressive inference using an MLP to determine if an accident has occurred and to provide a description of the surrounding environment.
Comparing the accuracy of accident detection with other video-based large language models, AccidentBlip2 stands out with an impressive accuracy of 66.5%, surpassing the performance of all baseline models. This highlights the effectiveness of our framework in complex multi-vehicle transportation systems. Specifically, our four-vehicle system achieves an accuracy of 72.2% in detecting environmental accidents, signifying a significant improvement over single-vehicle accident detection accuracy. Moreover, AccidentBlip2 demonstrates clear advantages in accident checking when compared to other video-only large language models used in multi-vehicle systems. These results validate the efficacy of our approach and its potential for enhancing accident detection capabilities in complex traffic scenarios.