### v4

# AccidentBlip2: Accident Detection With Multi-View MotionBlip2

Abstract
Intelligent vehicles have demonstrated excellent capabilities in many transportation scenarios, but the complex on-board sensors and the inference capabilities of on-board neural networks limit the accuracy of intelligent vehicles for accident detection in complex transportation systems. In this paper, we present AccidentBlip2, a pure vision-based multimodal large model Blip2 accident detection method. Our method first processes the multi-view through ViT-14g and inputs the multi-view features into the cross attention layer of the Qformer, while our self-designed Motion Qformer replaces the self-attention layer in Blip2’s Qformer with the Temporal Attention layer in the In the inference process, the query generated in the previous frame is input into the Temporal Attention layer to realize the inference for temporal information. Then we detect whether there is an accident in the surrounding environment by performing autoregressive inference on the query input to the MLP. We also extend our approach to a multi-vehicle cooperative system by deploying Motion Qformer on each vehicle and simultaneously inputting the inference-generated query into the MLP for autoregressive inference. Our approach detects the accuracy of existing video large language models and also adapts to multi-vehicle systems, making it more applicable to intelligent transportation scenarios.The code will be released at: https://github.com/YihuaJerry/AccidentBlip2.git
I INTRODUCTION
Traditional accident detection methods[1][2] have proven to be useful for single vehicles, their limitations become apparent when it comes to handling complex traffic scenarios and being deployed across multiple vehicles, which ultimately affects their reliability. Take, for instance, the utilization of LSTM and RNN for timing in traffic information detection—a method that can only be applied to a single viewpoint. Such an approach falls short when confronted with traffic scenarios that require multi-view detection, let alone the detection of accidents occurring in the environment involving multiple vehicles.
The utilization of Multimodal Large Language Models (MLLM) for accident detection[3] holds potential in addressing complex scenarios, its effectiveness is currently limited. The approach of employing MLLMs in conjunction with multi-sensor fusion proves impractical due to the high cost associated with training. Furthermore, relying solely on video-based MLLMs results in a notable decrease in accuracy when attempting joint multi-vehicle detection. The closest existing solution, Video-Llava, falls short in accurately detecting accidents within intricate environments and yields subpar results when handling multi-view inputs, primarily due to its reliance on feature mapping for video inference.
In this paper, we introduce AccidentBlip2, a Multimodal Language Model (MLLM) that utilizes multi-view Motion Qformer reasoning. Our objective is to enable the model to perform accurate multiview temporal inference. To achieve this, our approach incorporates ViT (Vision Transformer) to extract features from an image, which are then fused into the Qformer. Simultaneously, we input the queries generated during the Qformer inference process of the current frame into the Qformer corresponding to the next image, allowing it to participate in cross-attention computation. These queries possess the ability to inherit multiview information from the previous time sequence while simultaneously saving and updating the information of the current moment through the Qformer. Consequently, regardless of the scenario, the generated queries effectively preserve and transfer the time series information of each multi-view frame, thereby minimizing information loss between frames in the feature mapping process.
In addition to the perceptual analysis of the surrounding scene of a single vehicle, we have also developed a multi-vehicle collaborative end-to-end scenario perception scheme to compensate for some of the blind spots and deficiencies of single-vehicle perception. We extended the single-vehicle environmental experiments to end-to-end driving scenarios, testing the accuracy of accident judgment and perception in addition to self-vehicle and multi-vehicle association, respectively. Overall, our main contributions are as follows:
  • 1)
    We propose an innovative visual-only traffic accident determination agent that focuses on detecting the occurrence of accidents and alerting drivers to driving safety in complex driving environments.
  • 2)
    We present a framework for end-to-end accident detection based on MLLMs, enabling the LLMs to determine whether an accident has occurred or there is a an accident happened in or around a multi-vehicle system based on end-side vehicle information.
II RELATED WORK
II-A MultiModal Large Language Model
With the advent of GPT4, a large number of LLMs began to explore multimodal capabilities. These LLMs use multimodal information to fine-tune the language model to enhance the LLMs’ fusion ability and comprehension in different modalities. Among them, the visual-verbal MLLMs, led by Llava-v1.5[4], have spawned MLLMs applicable to different scenarios. In addition owlViT[5] and other MLLMs for target detection are fine-tuned for specific scenes by combining specific datasets, allowing them to accomplish specific visual tasks with user-prompt commands. For temporal multimodal information, researchers have also introduced Video-Llava[6], which incorporate the processing capabilities of temporal pictures and videos into a LLM.
II-B MLLMs for Automatic Driving
Recently, with the rapid development of autonomous driving models, LLMs are gradually appearing in the field of intelligent driving and complex traffic awareness. The most typical model is UniAD[7], which completely applies LLMs to the field of autonomous driving. UniAD integrates the three important parts of autonomous driving: perception, decision-making, and planning in a unified network architecture, effectively reducing the loss of information between different independent modules. In the perception task, DRIVEGPT4[8] uses YOLOv8 to detect common targets such as vehicles in each frame of the video and sends the obtained border coordinates to ChatGPT as linguistic information. However, since DriveGPT4 only does environment-aware tasks for a single vehicle, it cannot be practically deployed for end-to-end tasks with multi-vehicle linkage.
Figure 1: Architecture of MotionBlip2: We propose a multi-view temporal detection framework. We firstly input the initialized query $Q_{0}$ into the Qformer corresponding to the first frame of the multi-view images. , and at the same time input the features of the first frame of the multi-view processed by ViT into the first Qformer. A new query $Q_{1}$ will be generated by the Qformer inference and input into the next Qformer along with the features of the ring view map of the current frame processed by ViT. After completing the last frame of image processing, we input the generated query $Q_{n}$ into the MLP and perform autoregressive computation, if there is a risk of accident we will alert the driver and vice versa we will remind the driver to drive with caution.
II-C Accident Detection
Detection of traffic accidents as the most concerned research area in autonomous driving safety, many researchers have done a lot of work on accident detection. Based on traditional detect methods, the researchers used the front view of the vehicle in conjunction with a neural network for temporal detection, such as an LSTM or RNN, to warn drivers of a crash[1]. However, this method can only sense the vehicle travelling status around the bicycle, and cannot sense the danger in the complex traffic environment. However, due to the fact that its corpus is directly retrieved through LLMs, it is not able to reason about the perception of brand new environments in complex situations.
III METHODOLOGY
Blip2 currently lacks the capability to process 6-view picture inputs directly, as well as the ability to utilize temporal 6-view picture inference directly. To address this limitation, we leverage the CARLA[9] simulator’s simulation dataset and employ vision transforms to enhance Blip2’s[10] capacity in handling temporal 6-view pictures. Our objective is to train the Motion Qformer in such a way that it can effectively reason over temporal 6-view images, enabling them to be utilized for inference. It is worth noting that our framework can be extended to encompass multi-vehicle end-to-end autonomous driving simulation scenarios.
III-A Multi-view Input and Temporal Inference
We introduce a perceptual framework that leverages MLLMs, comprising of two key components: multi-view image processing and temporal inference. In the multi-view processing stage, we utilize ViT-14g to handle multiple views and extract the relevant image features. These features capture important visual information from different perspectives. Moving to the temporal inference stage, we employ queries as carriers of temporal information for forward transfer. These queries encapsulate contextual and temporal cues that facilitate understanding and reasoning over time. By combining these two stages, our framework enables comprehensive perception and inference in a multimodal manner.
Our visual encoder uses ViT-14g from EVA-CLIP[11]. Vision Transformer first reads the multiple-view images provided by the camera and extracts features separately after resizing each image. We use the pre-trained vision transformer as a backbone to extract the features $f_{t}$ the multi-view of $t$ frames, and input the features of each multi-view frame into the linear layer of the Qformer to do the inference. Before inputting the multi-view features of the $t$ frame into Qformer, the vision transformer will first spread the features of each image into one dimension, then stitch the features of the six images, and input the stitched features $f_{t}$ into Qformer’s cross-attention. At the same time, the learnable query $Q_{t}$ is input into only Qformer’s self-attention layer, and the new query is output through the feed forward layer.
Depending on the type of data in autopilot, we need to propose a framework that can handle temporal images. We propose temporal self-attention, which inputs the multi-view images’ information for each frame through the temporal query $Q_{n}$, enter $Q_{n}$ into the corresponding Qformer respectively, and obtain the query $Q_{n+1}$ corresponding to the input of the next frame. Taking advantages of attention mechanism in Qformer, it interacts with extracted features from last output of Qformer, as shown in 3. Each frame’s multi-view features $f_{t}$ present in attention with temporal query $Q_{n}$, written in 1,
$Attn(Q_{n-1},K_{Car},V_{Car})=Softmax(\frac{Q_{n}\cdot K_{Car}^{T}}{\sqrt{d_{k}}})V$ (1)
where $Q_{n-1}\in\mathbb{R}^{N_{Q}\times D}$ denotes the query feature from last timestamp, which incurs to generate the output of current state $Q_{n}$. After that, in order to present the current image view. the current state $Q_{n}$ combines the image features from ViT-g, denoted as $f_{t}$. This cross attention mechanism can be described as
$Q=Q_{n}W^{Q}\hskip 10.00002ptK=V=f_{t}W^{K}$ (2)
$CrossAttn(Q_{n},f_{t})=Softmax\left(\frac{QK^{T}}{\sqrt{d_{2}}}\right)V$ (3)
where $Q$ denotes the projection layer from the feature $Q_{n}$, $K$ and $V$ denotes the features from $f_{t}$.
III-B Multi-Vehicle End-to-End Sensing
To ensure reliable vehicle environment perception, relying solely on a single vehicle can be an inadequate approach due to the existence of blind zones in visual perception. To address this challenge, we have implemented a vision-only perception scheme for multi-vehicle interaction. In complex traffic systems, our vehicle neural network is designed not only to detect the accident involving the self-vehicle but also to detect the accidents involving other vehicles around the environment. This is achieved through a multi-vehicle joint training method. Consequently, we have successfully trained a multi-vehicle accident-aware agent using data from the DeepAccident dataset[3].
Figure 2: Architecture of Motion Qformer: We replace the self-attention layer in Qformer with temporal attention layer, and encode the temporal information by inputting the query $Q_{n-1}$ generated in the previous frame to the attention head of temporal attention. The encoded token is input into cross attention layer and reasoned with the multi-view images feature $f$ generated in the current frame to generate the query $Q_{n}$ corresponding to the current frame by feed forward layer.
In our multi-vehicle system, we employ a pre-trained ViT-14g model on each vehicle, denoted as $Car_{n}$, to extract the features $f$ from the current frame’s multi-view images of the self-vehicle. These features are then fed into our specially designed Motion Qformer module. Additionally, for the queries generated by AccidentBlip2, we combine them with the multi-vehicle queries and input them into the Motion Qformer in the subsequent time step. Finally, we transform this high-dimensional query into a one-dimensional tensor to be used as input to the MLP (Multi-Layer Perceptron), as illustrated in Equation 4.
$\mathbf{X_{i}}=\text{MLP}(\text{concat}(Q_{1},Q_{2},Q_{3},Q_{4},...\text{dim}=0))$ (4)
Figure 3: Architecture of End-to-end AccidentBlip2: Our proposed solution to address multi-vehicle collaborative incident awareness. We follow the model architecture of MotionBlip2 for each vehicle.
We input the token output from the MLP at frame $i$ into embedding, and the embedded token $X_{i}$ is subjected to autoregression, from which we determine whether there is an accident between multi-vehicle systems. Each query $Q_{n}$ denotes one perspective from one vehicle $Car_{n}$ in multi-vehicle system, concatenated with one and the other, aggregating to be 24 viewpoints for the coordination of features.
IV EXPERIMENT
In this section, we validate our system on the simulation dataset DeepAccident. We test it separately for single-vehicle perception to multi-vehicle perception. More precisely, we will focus on explaining the following:
  • 1)
    Can AccidentBlip2 determine accidents in complex traffic systems more efficiently than other vision-only solutions used in time-series systems?
  • 2)
    Can AccidentBlip2 sense accidents occurring around it more accurately than a single-vehicle system in a multi-vehicle end-to-end system?
IV-A Training Data and Evaluation Metric
For training and evaluation data, we have adopted a public open-source dataset, DeepAccident[3], taking advantanges of its full visual information upon vehicles and infrastructures. In the DeepAccident dataset, it consists of couples of scenarios that emulates those real-world potential car collisions, attached with 4 vehicles, 1 infrastructure and BEV point cloud annotations. Nevertheless, in contrast to those multimodal techniques, AccidentBlip2 is based on vision-exclusive paradigms, hence involving only visual images sampled from objects.
((a)) Hints for hazardous situations
((b)) Hints for the security situation
Figure 4: Agent for accident prediction and road condition alerts: In response to the temporal pictures, 4(a) is the agent alerts the driver if there is a potential hazard or impending hazard in the sequence, 4(b) is if there are no safety hazards in the sequence pictures, the agent alerts the driver to drive safely.
As proposed by DeepAccident[3], it consists of 12 types of accident scenarios that happens upon signalized and unsignalized intersections. The dataset has 57k annotated V2X frames, split by the ratio of 0.7, 0.15 and 0.15 for training, validation and test set respectively. Our training can be fit in a 4-GPU A6000 machine, which merely requires less than 24 hours for the result.
As for evaluation metrics, Accuracy between positive prediction of accidents happening and its opposite is measured by the given formula 5,
Acc=TP+TNNumber of $\text{Acc}=\frac{\text{TP}+\text{TN}}{\text{Number of samples}}$ (5)
where TP denotes the true positive samples where model predicts the accurate result, while TN stands for the true negative samples where model predicts as true sample but the scenario does not happen any accidents.
IV-B Implementation Details
Baselines. Our main baseline is the common Video LLMs (Video-Llama[12]). The use of Video-Llama is often seen when dealing with video tasks. we also compare it with Video-Vicuna, a fine-tuned version of Video-Llama. In addition to the Video-Llama family of video LLMs, we also used Video-Llava[6], a video inference large language model of alignment before projection, as a comparison. Purely visual fine-tuned LLMs like Llava[4] do not easily scale to temporal data inputs, and they are not superior to Video-Llama, so they are not included in the comparison.
Our proposed model consists of ViT-14g as the visual encoder, Motion Qformer leveraged as extracting and concatenating features from both texts and images. While taking into account the optimum experiment for accident prediction, we train AccidentBlip2 with image size of $224\times 224$.
In our setup, the parameters of our model are frozen for the visual model ViT-14g and the language model OPT-2.7B[14], which have been initialized with pre-trained weights, equivalent as Blip2[10]. Motion Qformer, as it sides with Qformer features, aims to fuse multiple view with collective vehicles together, loading and trained with tasks that supports for both prediction and language model outputs. Image input are passed through ViT layer into Motion Qformer in order to capture the temporal dynamic features within different frames.
During training, we selected a learning rate warm up for the initial 3 epochs, coupled with the Adam optimizer[15] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. The AccidentBlip2 model was trained for 8 epochs, using a cosine learning rate decay as the scheduler of learning rate up to $51e-$ 5. The training epoch is 6 and batch size is 8, varing from tasks involved with different numbers of vehicles. In terms of multiple scenarios, with sampling from timestamp 0 to T(final timestamp), the input shape is $X_{S}\in\mathbb{R}^{T\times N_{V}\times V_{C}\times C\times H\times W}$, while $T$ denotes the time length, $N_{V}$ represents the number of vehicle, $V_{C}$ denotes 6 separate view of the vehicle, encompassing Front, FrontLeft, FrontRight, Back, BackLeft and BackRight cameras, and $H$ and $W$ stand for the size of input image. Motion Qformer recursively interacts with the query output $Q_{T-1}\in\mathbb{R}^{N_{V}\times N_{Q}\times D}$, which spawns from the last timestamp viewpoints. It takes in and out the temporal multi-view features repetitively until reaching to its end, then LLM module is in charge of collecting the final Query Output $Q_{T}$ aggregating a comprehensive image embedding. Our training loss $Loss(p_{t})$, shown in following 6, which is calculated based on Focal Loss, where $p_{t}$ denotes the probability of true class, $\alpha$ is used to balance the weights of positive and negative samples and $\gamma$ take controls the rate of samples detected.
$Loss(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ (6)
Owing to mitigating imbalance categories as our model predicts on accidents, $\alpha$ is set to be 0.25, while $\gamma$ lies to its default value 2.0, conducive to train the model gaining greater focus towards misclassified samples.
Figure 5 shows the evaluation curve of our model on DeepAccident[2] dataset. Our models get a profound accuracy in different scenario, especially when it involves with the most amount of visual information of 4 vehicles.
Figure 5: Visualization of training process
IV-C Main Result
Different from the existing models, our AccidentBlip2 has a profound coordination in tackling with multiple vehicles and complex views in each of them. Instead of focusing on one viewpoint only, like Video-Llama AccidentBlip2 processes a number of temporal images simultaneously, while other video models are only able to input a large concatenated video, which reduces the precision in the recognition of accidents. Notwithstanding that V2XFormer collects all scenarios, involving BEV features and object location annotations, it incurs to bring redundancies mixed with unused information. Our introduced model leveraged vision-only frame images, outdoing in video scenarios than V2XFormer.
TABLE I: Performance Comparison Amongst Different Video Varge Language Models.
Model single vehicle(Acc.) multiple vehicles(Acc.)
Video-LLaMA 51.34 33.33
Video-Vicuna 36.43 31.87
Video-LLaVA 49.7 47.7
V2XFormer 43.8 69.5
AccidentBlip2 (ours) 66.5 73.1
Our proposed model outperform baseline models in both single-vehicle and multi-vehicle configurations. Compared with these video large language models, shown in table I, our AccidentBlip2 achieves 66.5% with merely single car input. In comparison with V2XFormer, proposed with three different V2X-agent fusion setting, a significant increment can be observed in our model, with about 3% improvement in accuracy.
Table II presents the results of performances of different different V2X configuration inputs in our model. By designing different V2X configuration models to demonstrating the model’s performance, we have evaluated AccidentBlip2 in single-car scenario, multi-car scenario, including ego vehicle, other vehicle and their cars behind, along with full-view scenario by introducing infrastructure view. Benefiting from the six camera sensors installed across four vehicles, our Motion Qformer effectively captures the temporal features of multiple vehicles.
TABLE II: Different V2X Configuration Models
Ego Vehicle Other Vehicle Vehicles Behind Infrastructure Acc.
$\checkmark$ -- -- -- 66.5
$\checkmark$ $\checkmark$ -- -- 68.5
$\checkmark$ $\checkmark$ -- $\checkmark$ 70.4
$\checkmark$ $\checkmark$ $\checkmark$ -- 73.1
This results in an accuracy improvement of 2% compared to the single-vehicle scenario. Moreover, an apparent increment by 6.6% can be seen with the advent of multiple views, which shows that the model effectively takes into account the overview of accidents.
Analysis on multi-vehicle paradigm. Our approach is designed to address the challenge of gradient shifts during training, which can lead to gradient explosion in the initial epochs. To mitigate this issue, we employ a warmup scheduler that makes slight adjustments to the learning rate within a minimal range. This helps stabilize the training process. Our multi-vehicle paradigm operates on four separate GPUs, allowing us to gather query results from each motion Qformer. Additionally, we conducted experiments to explore the impact of the initial weight of the Qformer, comparing pre-trained weights from Blip2 with default initialization. The results revealed a decrease in the overall performance of our language model when evaluated on the DeepAccident dataset. To overcome this, we utilize pre-trained weights from Blip2 to initialize AccidentBlip2, improving the model’s effectiveness in accident detection tasks.
V CONCLUSIONS
In this research paper, we propose a Motion Qformer-based accident detection framework named AccidentBlip2, which solely relies on vision inputs for analyzing road information. This framework introduces temporal attention in Blip2 by replacing self-attention mechanisms. It utilizes queries as carriers of temporal features, encoding the information from each frame into the temporal attention of the subsequent frame. This enables autoregressive inference using an MLP to determine if an accident has occurred and to provide a description of the surrounding environment.
Comparing the accuracy of accident detection with other video-based large language models, AccidentBlip2 stands out with an impressive accuracy of 66.5%, surpassing the performance of all baseline models. This highlights the effectiveness of our framework in complex multi-vehicle transportation systems. Specifically, our four-vehicle system achieves an accuracy of 72.2% in detecting environmental accidents, signifying a significant improvement over single-vehicle accident detection accuracy. Moreover, AccidentBlip2 demonstrates clear advantages in accident checking when compared to other video-only large language models used in multi-vehicle systems. These results validate the efficacy of our approach and its potential for enhancing accident detection capabilities in complex traffic scenarios.